### v1

@@ 2.4. Odometers
where $\tau_{n}\colon G/\Gamma_{n+1}\to G/\Gamma_{n}$ is the canonical projection, for every $n\in\mathbb{N}$. The space $X$ is a Cantor set when we endow every $G/\Gamma_{n}$ with the discrete topology, $\prod_{n\in\mathbb{N}}G/\Gamma_{n}$ with the product topology and $X$ with the induced topology. Let $\tau:G\to X$ given by $\tau(g)=(g\Gamma_{n})_{n\in\mathbb{N}}$. The group $G$ acts on $X$ as $g(g_{n}\Gamma_{n})_{n\in\mathbb{N}}=(gg_{n}\Gamma_{n})_{n\in\mathbb{N}}$, for every $g\in G$ and $(g_{n}\Gamma_{n})_{n\in\mathbb{N}}\in X$. The topological dynamical system $(X,G)$ is known as the $G$-odometer or odometer associated to the sequence $(\Gamma_{n})_{n\in\mathbb{N}}$. This system is equicontinuous and minimal. The $G$-odometer is free if and only if $\bigcap_{n\in\mathbb{N}}g_{n}\Gamma_{n}g_{n}^{-1}=\{1_{G}\}$, for every $(g_{n}\Gamma_{n})_{n\in\mathbb{N}}\in\overleftarrow{G}$, which implies that there exists a free $G$-odometer if and only if $G$ is residually finite (see [6]).
If the groups $\Gamma_{n}$’s are normal, we call the $G$-odometer an exact $G$-odometer or exact-odometer, and we denote it as $\overleftarrow{G}=\overleftarrow{G}_{(\Gamma_{n})}$. In this case, $\overleftarrow{G}$ is a subgroup of $\prod_{n\in\mathbb{N}}G/\Gamma_{n}$, and the function $\tau:G\to\overleftarrow{G}$ is a group homomorphism. Furthermore, the $G$ action corresponds to the coordinatewise left multiplication by $\tau(g)$, for every $g\in G$. On the other hand, $(\overleftarrow{G},G)$ is free if and only if $\bigcap_{n\in\mathbb{N}}\Gamma_{n}=\{1_{G}\}$. In this case, $\tau(G)$ is a dense subgroup of $\overleftarrow{G}$ isomorphic to $G$, and $\overleftarrow{G}$ is a totally disconnected metric compactification of $G$. Indeed, the totally disconnected metric compactifications of $G$ are precisely the free exact $G$-odometers (see [4, Lemma 2.1]).
@@
If $(X,G)$ is a minimal equicontinuous Cantor system, then $(X,G)$ is conjugate to a $G$-odometer.
@@
It is sufficient to follow the proof of [8, Theorem 2.7], and to note that if $(X,G)$ is not free, then the function $\varphi:G\to K$ is not necessarily injective. This implies that the action of $G$ on $K$ is not necessarily free. The intersection of the groups $G_{n}$’s mentioned there, is the stabilizer of the neutral element $e\in K$ (and maybe non-trivial). Then, the proof follows to conclude that $(X,G)$ is regularly recurrent, meaning that for every $x\in X$ and every clopen neighborhood $C$ of $x$, there exists a finite index subgroup $\Gamma$ of $G$ such that $\gamma x\in C$, for every $\gamma\in\Gamma$. Let $x\in X$ and $(U_{n})_{n\in\mathbb{N}}$ be a nested sequence of clopen sets such that $\bigcap_{n\in\mathbb{N}}U_{n}=\{x\}$. Then, by [6, Lemma 4], for every $n\in\mathbb{N}$ there exists a finite index subgroup $\Gamma_{n}$ such that $C_{n}=\overline{O_{\Gamma_{n}}(x)}\subseteq U_{n}$ is a clopen set verifying $C_{n}\cap\gamma C_{n}\neq\emptyset$ if and only if $\gamma C_{n}=C_{n}$, and if and only if $\gamma\in\Gamma_{n}$. The system $(X,G)$ is conjugate to the odometer given by the sequence $(\Gamma_{n})_{n\in\mathbb{N}}$ (see the proof of [6, Theorem 2]). ∎
The following result is a consequence of Proposition 2
@@
Let $G$ be a countable group. The following statements are equivalent:
@@
    There exists a minimal equicontinuous Cantor system $(X,G)$.
@@
    The group $G/R(G)$ is infinite, where $R(G)$ is the residual subgroup of $G$.
@@
    $G$ is an extension of an infinite residually finite group.
Proof.
Suppose that $(X,G)$ is a minimal equicontinuous system. By Proposition 2, the system $(X,G)$ is conjugate to a $G$-odometer. Let $(\Gamma_{n})_{n\in\mathbb{N}}$ be a decreasing sequence of finite index subgroups of $G$ such that $(X,G)$ is conjugate to the odometer associated to $(\Gamma_{n})_{n\in\mathbb{N}}$. Let $\Gamma=\bigcap_{n\in\mathbb{N}}\Gamma_{n}$. Since $X$ is a Cantor set (therefore, an infinite set), the groups $\Gamma_{n}$’s do not stabilize. This implies that $G/\Gamma$ is infinite. Since $R(G)\subseteq\Gamma$, we get that $G/R(G)$ is infinite. This shows (1) implies (2). Since $G/R(G)$ is residually finite, the implication (2) to (3) is direct from the exact sequence $1\to R(G)\to G\to G/R(G)\to 1$. Finally, if $G$ is an extension of an infinite residually finite group, then a normal subgroup $H$ of $G$ exists such that $G/H$ is infinite and residually finite. This implies there exists a strictly decreasing sequence $(\Gamma_{n})_{n\in\mathbb{N}}$ of finite index subgroup of $G/H$ such that $\bigcap\Gamma_{n}=\{1_{G/H}\}$. If $\pi:G\to G/H$ is the canonical projection, then $(\pi^{-1}(\Gamma_{n}))_{n\in\mathbb{N}}$ is a strictly decreasing sequence of finite index subgroups of $G$. The $G$-odometer associated to $(\pi^{-1}(\Gamma_{n}))_{n\in\mathbb{N}}$ is a minimal equicontinuous Cantor system given by the action of $G$. This shows (3) implies (1). ∎
2.5. Toeplitz subshifts
Let $\Sigma$ be a finite alphabet. An element $x\in\Sigma^{G}$ is Toeplitz if for every $g\in G$ there exists a finite index subgroup $\Gamma$ of $G$ such that $x(g)=x(\gamma^{-1}g)$, for every $\gamma\in\Gamma$. We say that $\Gamma$ is a period of $x$ or that the coordinate $g$ is periodic with respect to $\Gamma$. A subshift $X\subseteq\Sigma^{G}$ is a Toeplitz subshift if there exists a Toeplitz element $x\in\Sigma^{G}$, such that $X=\overline{O_{G}(x)}$. The Toeplitz subshifts are always minimal (see [6, 24, 25]). On the other hand, observe that if $x\in\Sigma^{G}$ is a non-periodic Toeplitz subshift (i.e., its stabilizer $\{g\in G:gx=x\}$ with respect to the shift action is trivial), then the intersection of its periods is trivial. This implies that $G$ is residually finite. In addition to the following proposition, this observation allows us to characterize residually finite groups as those that admit non-periodic Toeplitz elements.
@@ 2.5. Toeplitz subshifts
([6, Theorem 7]) Let $(\Gamma_{n})_{n\in\mathbb{N}}$ be a strictly decreasing sequence of finite index subgroups (not necessarily normal or with trivial intersection) of $G$. Then there exists a Toeplitz subshift $X\subseteq\{0,1\}^{G}$ such that $(X,G)$ is an almost 1-1 extension of the $G$-odometer associated to $(\Gamma_{n})_{n\in\mathbb{N}}$.
@@
([4, Lemma 2.4], [6]) Let $x\in\Sigma^{G}$ be a Toeplitz element. For every period $\Gamma$ of $x$, there exists an essential period $K$ of $x$ such that ${\rm Per}(x,\Gamma,\alpha)\subseteq{\rm Per}(x,K,\alpha)$, for every $\alpha\in\Sigma$.
A period structure of $x$ is a nested sequence $(\Gamma_{n})_{n\in\mathbb{N}}$ of essential periods of $x$, such that $G=\bigcup_{n\in\mathbb{N}}\Gamma_{n}$. A period structure of $x$ always exists (see [6, Corollary 6]). Let $(\Gamma_{n})_{n\in\mathbb{N}}$ be a period structure of $x$. Let $X=\overline{O_{G}(x)}$ be the associated Toeplitz subshift. For each $n\in\mathbb{N}$, consider the set
@@
([6, Lemma 8 and Proposition 6] or [4, Lemma 2.5]) For every $n\in\mathbb{N}$, the set $B_{n}$ is closed. In addition, for every $g,h\in G$ the following statements are equivalent:
@@
Then the collection $\{g^{-1}B_{n}:g\in D_{n}\}$ is a clopen partition of $X$, where $D_{n}$ is a subset of $G$ containing exactly one representative element of the classes in $G/\Gamma_{n}$, for every $n\in\mathbb{N}$.
Remark 7.
Observe that the stabilizer of $y\in X$ is the group $\bigcap_{n\in\mathbb{N}}v_{n}\Gamma_{n}v_{n}^{-1}$, where $v_{n}\in G$ is such that $y\in v_{n}B_{n}$. Thus, if the groups $\Gamma_{n}$ are normal subgroups of $G$, the Toeplitz subshift $(X,G)$ is free if and only if $\bigcap_{n\in\mathbb{N}}\Gamma_{n}=\{1_{G}\}$.
@@
The first part is direct from Lemma 6. The second part is a consequence of Proposition 2 and [25, Theorem 5.8]. ∎
3. Almost automorphic systems with no invariant measures.
Let $G$ be an infinite countable group, and let $(\Gamma_{n})_{n\in\mathbb{N}}$ be a strictly decreasing sequence of finite index subgroups of $G$. Observe that $\bigcap_{n\in\mathbb{N}}\Gamma_{n}$ is not of finite index. The following result is well known; we include it for the reader’s convenience (see, for example, [4, 7]).
Lemma 8.
Let $G$ be an infinite countable group, and let $(\Gamma_{n})_{n\in\mathbb{N}}$ be a strictly decreasing sequence of finite index subgroups of $G$. Let $H=\bigcap_{n\in\mathbb{N}}\Gamma_{n}$ and $D\subseteq G$ be a set that contains $1_{G}$ and exactly one element of each class $\{Hg:g\in G\}$. Then there exist an increasing sequence $(n_{i})_{i\in\mathbb{N}}\subseteq\mathbb{N}$ and a sequence $(D_{i})_{i\in\mathbb{N}}$ of finite subsets of $D$, such that for every $i\in\mathbb{N}$
@@
    $D_{i}$ contains exactly one element of each class $\{\Gamma_{n_{i}}g:g\in G\}$.
@@
    $D=\bigcup_{i\in\mathbb{N}}D_{i}$.
@@
Observe that if $S\subseteq D$ is a finite set, then there exists $l\geq 1$ such that if $g_{1}$ and $g_{2}$ are two different elements in $S$, then $\Gamma_{l}g_{1}\neq\Gamma_{l}g_{2}$. Indeed, suppose there exists a finite subset $S$ of $D$ such that for every $l\geq 1$ there are two different elements $g_{l}$ and $h_{l}$ in $S$ such that $\Gamma_{l}g_{l}=\Gamma_{l}h_{l}$. Since $S$ is finite, there exist $g,h\in S$ and a subsequence $(\Gamma_{l_{i}})_{i\in\mathbb{N}}$ such that $g_{l_{i}}=g$ and $h_{l_{i}}=h$, for every $i\in\mathbb{N}$. This implies that $hg^{-1}\in\bigcap_{i\in\mathbb{N}}\Gamma_{l_{i}}=\bigcap_{i\in\mathbb{N}}\Gamma_{i}=H$, which is a contradiction with the fact that $g\neq h\in S\subseteq D$.
Let $D=\{g_{1},g_{2},\dots\}$, with $g_{1}=1_{G}$. We will construct a sequence $(F_{i})_{i\in\mathbb{N}}$ of finite subsets of $D$ verifying points (1), (2) and (3): choose $F_{1}$ as a fundamental domain of $G/\Gamma_{1}$ such that $g_{1}\in F_{1}$. We can choose $F_{1}$ contained in $D$ by replacing the elements $g$ of $F_{1}$ which are not in $D$, by elements in $D$ which are in $Hg$. We set ${n_{1}}=1$. Let $S_{2}=F_{1}\cup\{g_{1},g_{2}\}\subseteq D$. By the previous discussion, there exists $\Gamma_{n_{2}}\subseteq\Gamma_{n_{1}}$ satisfying $\Gamma_{n_{2}}g\neq\Gamma_{n_{2}}h$ for every pair of distinct elements $g,h\in S_{2}$. Thus, we can take $F_{2}\subseteq D$, fundamental domain of $G/\Gamma_{n_{2}}$, such that $S_{2}\subseteq F_{2}$. Suppose we have defined $\Gamma_{n_{k-1}}$ and $F_{k-1}\subseteq D$ such that $\Gamma_{n_{k-1}}\subseteq\Gamma_{n_{k-2}}$ and $F_{k-1}$ is a fundamental domain of $\Gamma_{n_{k-1}}$ containing the set $F_{k-2}\cup\{g_{1},\dots,g_{k-1}\}$. Now, consider $S_{k}=\{g_{1},g_{2},\dots,g_{k}\}\cup F_{k-1}\subseteq D$. By the previous discussion, we can choose $\Gamma_{n_{k}}\subseteq\Gamma_{n_{k-1}}$ such that the elements in $S_{k}$ are in different equivalence classes of $G/\Gamma_{n_{k}}$. Thus, we can get $F_{k}\subseteq D$ satisfying $S_{k}\subseteq F_{k}$.
Thus, we obtain (1) and (2). Additionally,
$D=\bigcup_{i=1}^{\infty}S_{i}\subseteq\bigcup_{i=1}^{\infty}F_{i}\subseteq D,$
which implies (3).
Let $(Z,G)$ be the $G$-odometer associated to the strictly decreasing sequence $(\Gamma_{n})_{n\in\mathbb{N}}$ of finite index subgroups of $G$ (see Section 2.4). Let $D$ be a fundamental domain of $\bigcap_{n\in\mathbb{N}}\Gamma_{n}$ containing $1_{G}$. Taking subsequences if necessary, Lemma 8 ensures the existence of a sequence $(D_{i})_{i\in\mathbb{N}}$ of finite subsets of $D$ such that for every $i\in\mathbb{N}$
  1. (1)
    $1_{G}\in D_{1}$ and $D_{i}\subseteq D_{i+1}$.
  2. (2)
    $D_{i}$ is a fundamental domain of $\Gamma_{i}$.
  3. (3)
    $D=\bigcup_{i\in\mathbb{N}}D_{i}$.
For every $n\in\mathbb{N}$, we define
$C_{n}=\{(g_{i}\Gamma_{i})_{i\in\mathbb{N}}\in Z:g_{n}\Gamma_{n}=1_{G}\Gamma_{n}\}.$
It is straightforward to check that the topology of $Z$ is generated by the collection of sets $\{gC_{n}:g\in D_{n},n\in\mathbb{N}\}$ (see for example, [6]). Let $\nu$ be the unique invariant probability measure of $(Z,G)$. Observe that
$X_{0}=X_{0}(S)=\{(t_{i}\Gamma_{i})_{i\in\mathbb{N}}\in Z:G=\bigcup_{i\in\mathbb{N}}t_{n_{i}}\Gamma_{n_{i}}D_{n_{i-1}}\}.$
This set is invariant with respect to the $G$ action on $Y$. Indeed, if $(t_{i}\Gamma_{i})_{i\in\mathbb{N}}\in X_{0}$ then given $h,g\in G$, there exists $i\in\mathbb{N}$ such that $g^{-1}h\in t_{n_{i}}\Gamma_{n_{i}}D_{n_{i-1}}$. This implies that $h\in(gt_{n_{i}})\Gamma_{n_{i}}D_{n_{i-1}}$, from which we get that $g(t_{i}\Gamma_{i})_{i\in\mathbb{N}}\in X_{0}$.
Lemma 9.
There exists a subsequence $(\Gamma_{n_{i}})_{i\in\mathbb{N}}$ such that $\nu(X_{0})=0$, where
$X_{0}=\{(t_{i}\Gamma_{i})_{i\in\mathbb{N}}\in Z:G=\bigcup_{i\in\mathbb{N}}t_{n_{i}}\Gamma_{n_{i}}D_{n_{i-1}}\}.$
Proof.
Take $n_{0}=1$. Suppose we have chosen $n_{0}<\cdots<n_{i-1}$. We take $n_{i}>n_{i-1}$ such that
$\frac{|D_{n_{i-1}}|}{|D_{n_{i}}|}<\frac{1}{2^{i+1}}.$
Let $(t_{i}\Gamma_{i})_{i\in\mathbb{N}}\in Z$ and $g\in G$ be such that there exists $j\in\mathbb{N}$ for which $g\in t_{n_{j}}\Gamma_{n_{j}}D_{n_{j-1}}$. This implies that $g^{-1}t_{n_{j}}\in D_{n_{j-1}}^{-1}\Gamma_{n_{j}}$. From this, we get
$g^{-1}(t_{i}\Gamma_{i})_{i\in\mathbb{N}}\in\bigcup_{j\in\mathbb{N}}\bigcup_{d\in D_{n_{j-1}}^{-1}}dC_{n_{j}}.$
This implies
$X_{0}\subseteq\bigcap_{g\in G}\bigcup_{j\in\mathbb{N}}\bigcup_{d\in D_{n_{j-1}}^{-1}}gdC_{n_{j}}.$

### v2

@@ 2.3.1. Odometers
$=$ $\{(x_{n})_{n\in\mathbb{N}}\in\prod_{n\in\mathbb{N}}G/\Gamma_{n}:\tau_{n}(x_{n+1})=x_{n},\mbox{ for every }n\in\mathbb{N}\},$
where $\tau_{n}\colon G/\Gamma_{n+1}\to G/\Gamma_{n}$ is the canonical projection, for every $n\in\mathbb{N}$. The space $Z$ is a Cantor set when we endow every $G/\Gamma_{n}$ with the discrete topology, $\prod_{n\in\mathbb{N}}G/\Gamma_{n}$ with the product topology and $Z$ with the induced topology. Let $\tau:G\to Z$ be given by $\tau(g)=(g\Gamma_{n})_{n\in{\mathbb{N}}}$, for every $g\in G$. The group $G$ acts on $Z$ as $g(g_{n}\Gamma_{n})_{n\in{\mathbb{N}}}=(gg_{n}\Gamma_{n})_{n\in{\mathbb{N}}}$, for every $g\in G$ and $(g_{n}\Gamma_{n})_{n\in{\mathbb{N}}}\in Z$. The topological dynamical system $(Z,G)$ is known as the $G$-odometer or odometer associated to the sequence $(\Gamma_{n})_{n\in{\mathbb{N}}}$. This system is equicontinuous and minimal. The $G$-odometer is free if and only if $\bigcap_{n\in{\mathbb{N}}}g_{n}\Gamma_{n}g_{n}^{-1}=\{1_{G}\}$, for every $(g_{n}\Gamma_{n})_{n\in{\mathbb{N}}}\in Z$, which implies that there exists a free $G$-odometer if and only if $G$ is residually finite (see [4] for more details about odometers).
Let $(Z,G)$ be the $G$-odometer associated to the decreasing sequence $(\Gamma_{n})_{n\in{\mathbb{N}}}$ of finite index subgroups of $G$.
For every $n\in{\mathbb{N}}$, we define
$C_{n}=\{(g_{i}\Gamma_{i})_{i\in{\mathbb{N}}}\in Z:g_{n}\Gamma_{n}=1_{G}\Gamma_{n}\},$
and
${\mathcal{P}}_{n}=\{gC_{n}:g\in D^{-1}_{n},n\in{\mathbb{N}}\},$
where $D_{n}\subseteq G$ is a set containing exactly one element of each class in $\{\Gamma_{n}g:g\in G\}$. It is straightforward to check that the topology of $Z$ is generated by $({\mathcal{P}}_{n})_{n\in{\mathbb{N}}}$. Furthermore, $({\mathcal{P}}_{n})_{n\in{\mathbb{N}}}$ is a sequence of nested clopen partitions, and for every $n\in{\mathbb{N}}$, the intersection $gC_{n}\cap hC_{n}$ is non-empty if and only $g\in h\Gamma_{n}$, which is equivalent to $gC_{n}=hC_{n}$ (see, for example, [4]). We call the sets $gC_{n}$, basic clopen sets. Let $\nu$ be the unique invariant probability measure of $(Z,G)$. Observe that
Lemma 11.
A $G$-odometer has a free orbit if and only if it is conjugate to a $G$-odometer associated to a decreasing sequence of finite index subgroups of $G$ with trivial intersection.
Proof.
Let $Z$ be a $G$-odometer associated to the sequence $(\Gamma_{n})_{n\in{\mathbb{N}}}$ of finite index subgroups of $G$. Let $z=(z_{n}\Gamma_{n})_{n\in{\mathbb{N}}}\in Z$ be such that ${\rm Stab}(z)=\bigcap_{n\in{\mathbb{N}}}z_{n}\Gamma_{n}z_{n}^{-1}$ is trivial. For every $n\in{\mathbb{N}}$, let define $\tilde{\Gamma}_{n}=z_{n}\Gamma_{n}z_{n}^{-1}$. Let $Y$ be the $G$-odometer associated to the sequence $(\tilde{\Gamma}_{n})_{n\in{\mathbb{N}}}$. It is not difficult to show that the function $f:Y\to Z$, given by $f((g_{n}\tilde{\Gamma}_{n})_{n\in{\mathbb{N}}})=(g_{n}z_{n}\Gamma_{n})_{n\in{\mathbb{N}}}$ for every $(g_{n}\tilde{\Gamma}_{n})_{n\in{\mathbb{N}}}\in Y$, is a topological conjugacy. ∎
@@
$(X,G)$ is a minimal equicontinuous Cantor system if and only if $(X,G)$ is conjugate to a $G$-odometer.
@@
It is sufficient to follow the proof of [6, Theorem 2.7], and to note that if $(X,G)$ is not free, then the function $\varphi:G\to K$ is not necessarily injective. This implies that the action of $G$ on $K$ is not necessarily free. The intersection of the groups $G_{n}$’s mentioned there, is the stabilizer of the neutral element $e\in K$ (and maybe non-trivial). Then, the proof follows to conclude that $(X,G)$ is regularly recurrent, meaning that for every $x\in X$ and every clopen neighborhood $C$ of $x$, there exists a finite index subgroup $\Gamma$ of $G$ such that $\gamma x\in C$, for every $\gamma\in\Gamma$. Let $x\in X$ and $(U_{n})_{n\in{\mathbb{N}}}$ be a nested sequence of clopen sets such that $\bigcap_{n\in{\mathbb{N}}}U_{n}=\{x\}$. Then, by [4, Lemma 4], for every $n\in{\mathbb{N}}$ there exists a finite index subgroup $\Gamma_{n}$ such that $C_{n}=\overline{O_{\Gamma_{n}}(x)}\subseteq U_{n}$ is a clopen set verifying $C_{n}\cap\gamma C_{n}\neq\emptyset$ if and only if $\gamma C_{n}=C_{n}$, and if and only if $\gamma\in\Gamma_{n}$. The system $(X,G)$ is conjugate to the odometer given by the sequence $(\Gamma_{n})_{n\in{\mathbb{N}}}$ (see the proof of [4, Theorem 2]). ∎
The following result is a direct consequence of Lemma 11 and Proposition 12.
@@
Let $G$ be an infinite countable group. The following statements are equivalent:
@@
    $G$ is residually finite
@@
    There exists a free $G$-odometer.
@@
    There exists a free minimal equicontinuous Cantor system $(X,G)$.
  4. (4)
    There exists a minimal equicontinuous Cantor system $(X,G)$ with a free orbit.
2.3.2. Toeplitz subshifts
Let $\Sigma$ be a finite alphabet. An element $x\in\Sigma^{G}$ is Toeplitz if for every $g\in G$ there exists a finite index subgroup $\Gamma$ of $G$ such that $x(g)=x(\gamma^{-1}g)$, for every $\gamma\in\Gamma$. We say that $\Gamma$ is a period of $x$ or that the coordinate $g$ is periodic with respect to $\Gamma$. A subshift $X\subseteq\Sigma^{G}$ is a Toeplitz subshift if there exists a Toeplitz element $x\in\Sigma^{G}$, such that $X=\overline{O_{G}(x)}$. The Toeplitz subshifts are always minimal (see [4, 9, 16, 17] for more details on Toeplitz subshifts). Observe that if $x\in\Sigma^{G}$ is a non-periodic Toeplitz subshift (i.e., its stabilizer with respect to the shift action is trivial), then the intersection of its periods is trivial. From this observation we deduce that the existence of non-periodic Toeplitz elements in $\Sigma^{G}$ implies that $G$ is residually finite. The following proposition tell us the converse holds as well.
@@ 2.3.2. Toeplitz subshifts
([4, Theorem 7]) Let $(\Gamma_{n})_{n\in{\mathbb{N}}}$ be a strictly decreasing sequence of finite index subgroups (not necessarily normal or with trivial intersection) of $G$. Then there exists a Toeplitz subshift $X\subseteq\{0,1\}^{G}$ such that $(X,G)$ is an almost 1-1 extension of the $G$-odometer associated to $(\Gamma_{n})_{n\in{\mathbb{N}}}$.
@@
([3, Lemma 2.4], [4]) Let $x\in\Sigma^{G}$ be a Toeplitz element. For every period $\Gamma$ of $x$, there exists an essential period $K$ of $x$ such that ${\rm Per}(x,\Gamma,\alpha)\subseteq{\rm Per}(x,K,\alpha)$, for every $\alpha\in\Sigma$.
A period structure of $x$ is a nested sequence $(\Gamma_{n})_{n\in\mathbb{N}}$ of essential periods of $x$, such that $G=\bigcup_{n\in\mathbb{N}}\Gamma_{n}$. A period structure of $x$ always exists (see [4, Corollary 6]). Let $(\Gamma_{n})_{n\in\mathbb{N}}$ be a period structure of $x$. Let $X=\overline{O_{G}(x)}$ be the associated Toeplitz subshift. For each $n\in\mathbb{N}$, consider the set
@@
([4, Lemma 8 and Proposition 6] or [3, Lemma 2.5]) For every $n\in\mathbb{N}$, the set $B_{n}$ is closed. In addition, for every $g,h\in G$ the following statements are equivalent:
@@
Then the collection $\{g^{-1}B_{n}:g\in D_{n}^{-1}\}$ is a clopen partition of $X$, where $D_{n}$ is a subset of $G$ containing exactly one representative element of the classes $\{\Gamma_{n}g:g\in G\}$, for every $n\in{\mathbb{N}}$.
@@
The first part is direct from Lemma 16. The second part is a consequence of Proposition 12 and [17, Theorem 5.8]. ∎
3. Subsequences of finite index subgroups.
Let $G$ be an infinite residually finite group. Let $(Z,G)$ be a $G$-odometer associated to a sequence $(\Gamma_{n})_{n\in{\mathbb{N}}}$ of finite index subgroups of $G$. For every $n\in{\mathbb{N}}$, let $C_{n}$ be the clopen set and ${\mathcal{P}}_{n}$ be the partition introduced in Section 2.3.1. Let $(Y,G)$ be a minimal topological dynamical system such that there exists a factor map $\pi:Y\to Z$. For every $y\in Y$ and $n\in{\mathbb{N}}$, we define $t_{n}=t_{n,y}$ as the element in $D_{n}^{-1}$ such that
$\pi(y)\in t_{n,y}C_{n}.$
Observe that $\pi(y)=(t_{n,y}\Gamma_{n})_{n\in{\mathbb{N}}}$. Thus, for every $g\in G$ and $n\in{\mathbb{N}}$, we have
(3.1) $gt_{n,y}=t_{n,gy}\gamma,\mbox{ for some }\gamma\in\Gamma_{n}.$
The goal of this section is to show the following result:
Proposition 18.
Let $G$ be an infinite residually finite group, and let $(Z,G)$ be a minimal equicontinuous topological dynamical system with a free orbit. Let $(Y,G)$ be a minimal topological dynamical system such that there exists a factor map $\pi:Y\to Z$. Then there exists a sequence $(\Gamma_{n})_{n\in{\mathbb{N}}}$ of finite index subgroups of $G$ such that $(Z,G)$ is conjugate to the $G$-odometer associated to $(\Gamma_{n})_{n\in{\mathbb{N}}}$ and such that there exists an increasing sequence $(D_{n})_{n\in\mathbb{N}}$ of finite subsets of $G$, such that
  • (P1)
    $1_{G}\in D_{1}$ and $D_{n}\subseteq D_{n+1}$, for every $n\in\mathbb{N}$;
  • (P2)
    $D_{n}$ contains exactly one element of each class $\{\Gamma_{n}g:g\in G\}$, for every $n\in\mathbb{N}$;
  • (P3)
    $G=\bigcup_{n\in\mathbb{N}}D_{n}$;
  • (P4)
    $D_{n+1}=\bigcup_{v\in D_{n+1}\cap\Gamma_{j}}vD_{j}$, for every $j\leq n$ and $n\in\mathbb{N}$;
  • (P5)
    $\nu(Z_{0})=0$, where
    $Z_{0}=\{(t_{n}\Gamma_{n})_{n\in{\mathbb{N}}}\in Z:G=\bigcup_{n\in{\mathbb{N}}}t_{n}\Gamma_{n}D_{n-1}\}$
    and $\nu$ is the unique invariant probability measure of $(Z,G)$;
  • (P6)
    For every $y\in Y$ and $n\in{\mathbb{N}}$ there exists $h_{n}\in D_{n+1}\cap\Gamma_{n}t_{n,y}^{-1}$ such that
    $h_{n}\in h_{n}D_{n}\subseteq D_{n+1}\setminus D_{n},$
    and
    $\lim_{n\to\infty}h_{n}^{-1}y_{0}=y,$
    where $y_{0}$ is a fixed element in $\bigcap_{n\in{\mathbb{N}}}\pi^{-1}C_{n}$;
  • (P7)
    $\frac{|D_{n-1}|}{|D_{n}|}<\frac{1}{2^{n+1}}$, for every $n\in{\mathbb{N}}$.
We need to prove some lemmas in order to show Proposition 18. With this purpose, in the rest of this section, $(Z,G)$ and $(Y,G)$ are the dynamical systems given in Proposition 18. According to Lemma 11 and Proposition 12, the system $(Z,G)$ is conjugate to the $G$-odometer associated to a decreasing sequence $(\Gamma_{n})_{n\in{\mathbb{N}}}$ of finite index subgroups of $G$ with trivial intersection.
It is not difficult to check that the hypothesis of normality of the finite index subgroups in [5, Lemma 3] and [3, Lemma 2.8] can be avoided. Thus
we can assume that $(\Gamma_{n})_{n\in{\mathbb{N}}}$ is such that there exists a sequence $(D_{i})_{i\in\mathbb{N}}$ of finite subsets of $G$ such that for every $i\in\mathbb{N}$
@@
    $D_{i}$ contains exactly one element of each class $\{\Gamma_{i}g:g\in G\}$.
@@
    $G=\bigcup_{i\in\mathbb{N}}D_{i}$.
  4. (4)
    $D_{i+1}=\bigcup_{v\in D_{i+1}\cap\Gamma_{j}}vD_{j}$, for every $j\leq i$.
$Z_{0}(S)=\{(t_{i}\Gamma_{i})_{i\in{\mathbb{N}}}\in Z:G=\bigcup_{i\in{\mathbb{N}}}t_{n_{i}}\Gamma_{n_{i}}D_{n_{i-1}}\}.$
This set is invariant with respect to the $G$ action on $Z$. Indeed, if $(t_{i}\Gamma_{i})_{i\in{\mathbb{N}}}\in Z_{0}(S)$ then given $h,g\in G$, there exists $i\in{\mathbb{N}}$ such that $g^{-1}h\in t_{n_{i}}\Gamma_{n_{i}}D_{n_{i-1}}$. This implies that $h\in(gt_{n_{i}})\Gamma_{n_{i}}D_{n_{i-1}}$, from which we get that $g(t_{i}\Gamma_{i})_{i\in{\mathbb{N}}}\in Z_{0}(S)$.
Lemma 19.
If $S=(n_{i})_{i\in{\mathbb{N}}}$ is an increasing sequence of positive integers such that
$\frac{|D_{n_{i-1}}|}{|D_{n_{i}}|}<\frac{1}{2^{i+1}},\mbox{ for every }i\in{\mathbb{N}},$
then $\nu(Z_{0}(S))=0$.
@@
Observe that $z=(t_{i}\Gamma_{i})_{i\in{\mathbb{N}}}\in Z_{0}(S)$ if and only if for every $g\in G$ there exists $i\in{\mathbb{N}}$ such that $t_{n_{i}}^{-1}g\in\Gamma_{n_{i}}D_{n_{i}-1}$. This is equivalent to $g^{-1}t_{n_{i}}=v\gamma$, for some $\gamma\in\Gamma_{n_{i}}$ and $v\in D_{n_{i}-1}^{-1}$. And the previous relation is satisfied if and only if $g^{-1}z\in\bigcup_{d\in D_{n_{i-1}}^{-1}}dC_{n_{i}}$. This shows
(3.2) $\bigcap_{g\in G}\bigcup_{j\in{\mathbb{N}}}\bigcup_{d\in D_{n_{j-1}}^{-1}}gdC_{n_{j}}=Z_{0}(S).$
The relation (3.2) implies that $Z_{0}(S)$ is an invariant measurable set.